OpenStudy (leozap1):

*Fan & Medal* Please explain how I can do this, I want to learn it. Let f(x) = -2x + 4 and g(x) = -6x - 7. Find f(x) - g(x)

OpenStudy (jango_in_dtown):

HI... I can help you

OpenStudy (leozap1):

Okay... How do I solve this?

OpenStudy (jango_in_dtown):

Hmm... you should find f(x)-g(x)=(-2x+4)-(-6x-7)=-2x+6x+4+7=4x-11

OpenStudy (jango_in_dtown):

\[ax-bx=(a-b)x\]

OpenStudy (jango_in_dtown):

Just using this formula we did the sum here

OpenStudy (jango_in_dtown):

@leozap1 did you get it?

OpenStudy (leozap1):

I'm trying just a sec

OpenStudy (leozap1):

On this part "2x+6x+4+7" aren't we suppose to subtract? It says to find f(x) - g(x)

OpenStudy (jango_in_dtown):

yeah correct

OpenStudy (jango_in_dtown):

-(-6)=6

OpenStudy (leozap1):

So it would be -2x-6x-4-7?

OpenStudy (jango_in_dtown):

no

OpenStudy (jango_in_dtown):

wait

OpenStudy (leozap1):

Ohh so since they are already negative, it makes them positive?

OpenStudy (jango_in_dtown):

the expression of f(x) will not be changed. but the expression of g(x) will be its negative.. g(x)=-6x-7 so -g(x)=-(-6x-7)=6x+7

OpenStudy (jango_in_dtown):